OpenStudy (anonymous):

The midpoint of line segment from p1 p2 is (-1,4)If p1=(-3,4) What is p2?

OpenStudy (anonymous):

Let, P2(x, y), and given that P1(-3, 4) then, (x+3)/2 = -3, ===> x+3 = -6, x = -6-3 = -9, Also, (y-4)/2 = 1, ===> y-4 = 2, y = 2+4 = 6 Hence, P2 is : (-9, 6) >=========================< ANSWER